# Recursions, Trains, Trees, and Combinatorial Rod Set Algebra

### v1

@@ Introduction
Soon we will use our rod model to study this observation about the Fibonacci sequence by converting trains composed of rods of lengths $1$ and $2$ into trains composed of rods of lengths $1$, 33, and $4$. When a 22 is followed by a $1$ we can glue them together to make a $3$, and when 22 is followed by a $2$ we glue to make a $4$. This construction may end with an extra rod of length $2$ at the end of the train. Figure 2 shows an example.
@@ 0 $\Lambda$ 1 1
  • $\operatorname{F}(n,[2,3])=1,0,1,1,1,2,2,3,4,5,7,9,12,\ldots$, a Padovan sequence [14].
@@
The number of rods of each length is the Catalan number with index shifted by $1$ from the train counts. The recursion corresponds to the well-known convolution that the Catalan’s satisfy.
@@ 2 Trees
Here we explore a natural way to organize $\operatorname{Trains}(R)$ in a tree111 Sorry for the awkward metaphor mixing trees and trains. that lets us sytematically study the way we combined rods in Figure 2.
Definition 8.
@@ 3 Discrepancies
Let $A$ be the subset of $\operatorname{Trains}(R)$ for which $m=0$ in Lemma 11, and let $B=\operatorname{Trains}(R)\setminus A$.
@@ 5 Rod Set Algebra
If all the rods $Q$ we want to expand are already in $R$ then the rod set $|\text{Leaves}(\mathfrak{T})|$ for the expansion subtree $\mathfrak{T}$ of $\operatorname{Tree}(R)$ with $|\text{Inner}(\mathfrak{T})|=Q$ is $(R\setminus Q)\cup QR$. Theorem 15 then implies
@@
We use any convenient expansion subtree $\operatorname{Tree}(R^{\prime})$ for a rod set equivalent to $R$ and invoke Lemma 11 and Theorem 12. ∎
@@
Theorem 18 shows that there is a simple formula for $S$ in terms of $R$ and $Q$. The next theorem proves that given any two of $R,Q,S$, the third is well-defined.
Theorem 23.
@@
Theorem 18 tells us how to find $S$ given $R$ and $Q$.
Given $R$ and $S$ we can use Theorem 21 to find $Q$ by computing the discrepancies $\operatorname{D}(n,R,S)$.
@@
For finite $R$ and $S$ the $Q$ in $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ need not be finite. For example, when we use Algorithm 24 to find the $Q$ that mediates the expansion of the Fibonacci rod set $[1,2]$ to the Padovan rod set $[2,3]$ we expand the $1$, add a $3/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}$ pair, then expand the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} forever. The result is
@@
then Theorem 7 implies there is a rod set $R$ whose net train counts are $1,\mathscr{L}$. But that rod set is uninformative: it does not clearly connect the Lucas number sequence to the Fibonacci recursion determined by the rod set $[1,2]$. However, Theorem 21 says that the rod set $Q$ for which
@@
and $Q=[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$. Thus, also by Theorem 21, the Lucas numbers satisfy the recursion
@@
Theorem 21 provides more. Equation 5.3 shows a nice way to view the trains counted by $\mathscr{L}$:
@@
The argument in this Lucas number example relies only on Theorem 21, not on the finiteness of $S=[1,2]$ nor on the fact that the discrepancies occur only for values $n\leq\max S$. But when these finiteness conditions are met the result shows how to use our rod model to define “traditional” initial conditions.
@@ 6 Generating functions
If $R$ expands to $S$ via $Q$ then Theorem 18 implies $S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ so
@@
The Exchange Corollary 19 is an immediate consequence of Theorem 30. The Odd Sign Swap Corollary 20 follows when we replace $x$ with x-x in Equation 6.3.
This theorem shows that the Expansion Theorem 18 is essentially equivalent to the standard algorithm for multiplying formal power series. The net count of the rods of length $n$ in $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ encodes the calculation of the coefficient of $x^{n}$ in the product power series. Then the Composition Corollary 22 is just the distributive property. Algorithm 24 illustrated in Figure 6 calculates the quotient of two power series.
@@
The next corollary is a reprise of Theorem 23.
Corollary 32.
@@
The second assertion follows from Theorem 30. ∎
@@
The next corollary is a reprise of Theorem 26.
Corollary 34.
@@
The following theorem is a reprise of Theorem 7.
Theorem 35.
@@
We end this section with a well-known theorem on rational generating functions[7]:
Theorem 36.
@@
Since $Q$ specifies the discrepancies in this expansion, Theorem 25 implies it is finite, so
@@ 7 Duality and its Applications
The union of the internal nodes is $\operatorname{Trains}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})$. The leaves on the left are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k}R$ starting at $k=0$, while the leaves on the right are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k-1}QR$, starting at $k=1$. Equation 15 shows that all but the leaves in $R$ cancel. ∎
Corollary 39.
@@
If $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $S\xrightarrow{\hphantom{a}Q^{*}\hphantom{a}}R$ so $R\xrightarrow{\hphantom{a}Q^{**}\hphantom{a}}S$. Then Theorem 18 implies $Q^{**}\equiv Q$. ∎
@@
When $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, Theorem 18 shows how to compute $S$ given $R$ and $Q$, and Theorem 23 tells us that any two of $R,Q,S$ determines the third. With duality, it is straightforward, albeit a bit messy, to find explicit rod set formulas for $Q$ in terms of $R$ and $S$ and for $R$ in terms of $Q$ and $S$.
@@
For example, recall that we established the Fibonacci identities in Equations 3.2 and 3.3 using the third expansion tree in Figure 4 together with Theorem 21. We started with the infinite expansion
@@
Now we use Theorem 38 to dualize Equation 7.1:
@@
Theorem 12 implies the train counts for the rod set $1,3,5,7,\ldots]$ satisfy the Fibonacci recursion when $n>2$. The train counts for the rod set of odd numbers are the number of compositions, $\mathscr{O}(n)$, of $n$ into odd parts. Since $\mathscr{O}(1)=\mathscr{O}(2)=1$, the sequence $\mathscr{O}(n)$ is the Fibonacci sequence with these initial conditions, hence $\mathscr{O}(n)=\operatorname{F}(n-1,[1,2])$. For example, there are $\operatorname{F}(4,[1,2])=5$ rod trains of length $5$ built from odd length rods: $5,311,131,113,11111$.
@@
Since $R=[2,2,3]$ is an expansion of the Fibonacci rod set $[1,2]$ by the singleton rod set $[1]$, Corollary 22 implies $1,2]\xrightarrow{\hphantom{a}\hphantom{a}}S$ has discrepancies $3,6,9,\ldots]\cup[1,4,7,\ldots]$. The discrepancies are the rods with lengths $n\equiv 0,1\pmod{3}$, which leads to the following identities for the Fibonacci sequence $\operatorname{F}(n)=\operatorname{F}(n,[1,2])$:
@@ 8 Rod Sets and Binomial Coefficients
The two ways of writing this sum correspond to the dark diagonals in Figure 8. The first sum counts rods of length $b$ and the second rods of length $a$. Above the dark diagonal, the diagonal for $\operatorname{F}(n,[a,b])=0$ is empty (dashed in the figure) unless $n$ is a multiple of $a$ or $b$. Below the dark diagonal the train counts are positive. In practice, it’s hard to draw the diagonals for large $a$. Hopkins [8, Appendix A] shows similar diagrams.
@@
Theorem 42 leads to some nice binomial identities. The train counts for $R=[1,1]$ are powers of two. The corresponding “diagonals” are the rows of the triangle, providing another proof of the familiar identity
@@ 9 Periodicity
Theorem 25 says that that when $n>4$, the net train counts $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$ satisfy the recursion $\operatorname{F}(n)=\operatorname{F}(n-6)$ determined by rod set $[6]$.
@@
This is the special case of Theorem 25 when $S$ contains just a single positive rod. ∎
@@
Since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] to $[2k]$, Corollary 22 implies $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$. ∎
@@
Theorem 44 says periodicity is equivalent to $R$ expanding to the one element rod set $[p]$, for which $\operatorname{\mathcal{C}}(x,[p])=x^{p}$. When $R$ is finite Theorem 30 then implies (1) and (5) are equivalent.
@@
We encourage the reader to check this result using Algorithm 24. Theorem 45 says we can stop when we’ve expanded to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{15}$}.
@@ 10 Expandability and Scaling
Since $\max Q=2$, Theorem 25, tells us that the $\operatorname{D}(n,R,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{4}])$ must vanish for $n\geq 3$. In particular, $\operatorname{F}(3)+4F(3-4)=\operatorname{F}(3)=0$ and $\operatorname{F}(4)+4F(0)=-4+4=0$.
@@
Theorem 25 implies $a=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>\max Q=a-\max R$. Thus for positive $i<\max R$
@@
To prove the converse, suppose $\operatorname{F}(a-i,R)=0$ for $i=1,2,\ldots,\max R-1$. Let $Q$ be the rod set that expands $R$ to the rod set $S$ with $\operatorname{F}(a,R)$ copies of $a$, with $a$ an antirod if $\operatorname{F}(a,R)$ is negative. The assumption on the net train counts of $R$ say that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $a-\max R<n<a$. Setting $\operatorname{C}(a,S)=\operatorname{F}(a,R)$ guarantees discrepancy 0 at $n=a$. Then Corollary 27 implies $Q$ is finite. ∎
@@
In Theorem 50 we saw that 1-expansions of $R$ correspond to sequences of 0’s in $\operatorname{F}(n,R)$. For 2-expandability the correspondence is to divisibility among net train counts. For example, we will see soon that the expansion above of $[1,2]$ to $8^{48},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{7}]$ is equivalent to the fact that
@@
The Exchange Corollary 19 implies
@@
Suppose there is a finite expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ with $\operatorname{Shape}(S)=\langle a,b\rangle$. Theorem 25 implies $b=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>b-\max R$. In the range $b-\max R<n<b$, we have $\operatorname{F}(n-b,R)=0$.
@@
Let $Q$ be the rod set that expands $R$ to the $S$ with shape $\langle a,b\rangle$ and those net rod counts. It follows that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $b-\max R<n\leq b$. Corollary 27 implies $Q$ is finite. ∎

### v2

@@ Abstract
keywords:
recursion, compositions, Fibonacci, Padovan, generating function, Lucas sequence, Borwein polynomial
@@ Introduction
Soon we will use our rod model to study this observation about the Fibonacci sequence by converting trains composed of rods of lengths $1$ and $2$ into trains composed of rods of lengths $1$, 33, and $4$. When a $2$ is followed by a $1$ we can glue them together to make a $3$, and when $2$ is followed by a $2$ we glue to make a $4$. This construction may end with an extra rod of length $2$ at the end of the train. Figure 2 shows an example.
@@ 0 $\Lambda$ 1 1
  • $\operatorname{F}(n,[2,3])=1,0,1,1,1,2,2,3,4,5,7,9,12,\ldots$, a Padovan sequence [8, p 89] [16].
@@
The number of rods of each length is the Catalan number with index shifted by $1$ from the train counts. The recursion corresponds to the well-known convolution that the Catalans satisfy.
@@ 2 Trees
Here we explore a natural way to organize $\operatorname{Trains}(R)$ in a tree [^1] Sorry for the awkward metaphor mixing trees and trains. that lets us sytematically study the way we combined rods in Figure 2.
Definition 2.1.
@@ 3 Discrepancies
Let $A$ be the subset of $\operatorname{Trains}(R)$ for which $m=0$ in Lemma 3.2, and let $B=\operatorname{Trains}(R)\setminus A$.
@@ 5 Rod Set Algebra
If all the rods $Q$ we want to expand are already in $R$ then the rod set $|\text{Leaves}(\mathfrak{T})|$ for the expansion subtree $\mathfrak{T}$ of $\operatorname{Tree}(R)$ with $|\text{Inner}(\mathfrak{T})|=Q$ is $(R\setminus Q)\cup QR$. Theorem 4.3 then implies
@@
We use any convenient expansion subtree $\operatorname{Tree}(R^{\prime})$ for a rod set equivalent to $R$ and invoke Lemma 3.2 and Theorem 3.3. ∎
@@
Theorem 5.2 shows that there is a simple formula for $S$ in terms of $R$ and $Q$. The next theorem proves that given any two of $R,Q,S$, the third is well-defined.
Theorem 5.7.
@@
Theorem 5.2 tells us how to find $S$ given $R$ and $Q$.
Given $R$ and $S$ we can use Theorem 5.5 to find $Q$ by computing the discrepancies $\operatorname{D}(n,R,S)$.
@@
For finite $R$ and $S$ the $Q$ in $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ need not be finite. For example, when we use Algorithm 5.8 to find the $Q$ that mediates the expansion of the Fibonacci rod set $[1,2]$ to the Padovan rod set $[2,3]$ we expand the $1$, add a $3/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}}$ pair, then expand the $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{3}$} forever. The result is
@@
then Theorem 1.7 implies there is a rod set $R$ whose net train counts are $1,\mathscr{L}$. But that rod set is uninformative: it does not clearly connect the Lucas number sequence to the Fibonacci recursion determined by the rod set $[1,2]$. However, Theorem 5.5 says that the rod set $Q$ for which
@@
and $Q=[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}]$. Thus, also by Theorem 5.5, the Lucas numbers satisfy the recursion
@@
Theorem 5.5 provides more. Equation 5.3 shows a nice way to view the trains counted by $\mathscr{L}$:
@@
The argument in this Lucas number example relies only on Theorem 5.5, not on the finiteness of $S=[1,2]$ nor on the fact that the discrepancies occur only for values $n\leq\max S$. But when these finiteness conditions are met the result shows how to use our rod model to define “traditional” initial conditions.
@@ 6 Generating functions
If $R$ expands to $S$ via $Q$ then Theorem 5.2 implies $S\equiv R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ so
@@
The Exchange Corollary 5.3 is an immediate consequence of Theorem 6.3. The Odd Sign Swap Corollary 5.4 follows when we replace $x$ with x-x in Equation 6.3.
This theorem shows that the Expansion Theorem 5.2 is essentially equivalent to the standard algorithm for multiplying formal power series. The net count of the rods of length $n$ in $R\cup{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}\cup QR$ encodes the calculation of the coefficient of $x^{n}$ in the product power series. Then the Composition Corollary 5.6 is just the distributive property. Algorithm 5.8 illustrated in Figure 6 calculates the quotient of two power series.
@@
The next corollary is a reprise of Theorem 5.7.
Corollary 6.5.
@@
The second assertion follows from Theorem 6.3. ∎
@@
The next corollary is a reprise of Theorem 5.10.
Corollary 6.7.
@@
The following theorem is a reprise of Theorem 1.7.
Theorem 6.8.
@@
We end this section with a well-known theorem on rational generating functions[17, p 202]:
Theorem 6.9.
@@
Since $Q$ specifies the discrepancies in this expansion, Theorem 5.9 implies it is finite, so
@@ 7 Duality and its Applications
The union of the internal nodes is $\operatorname{Trains}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}})$. The leaves on the left are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k}R$ starting at $k=0$, while the leaves on the right are ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{Q}}^{k-1}QR$, starting at $k=1$. Equation 4.3 shows that all but the leaves in $R$ cancel. ∎
Corollary 7.3.
@@
If $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, then $S\xrightarrow{\hphantom{a}Q^{*}\hphantom{a}}R$ so $R\xrightarrow{\hphantom{a}Q^{**}\hphantom{a}}S$. Then Theorem 5.2 implies $Q^{**}\equiv Q$. ∎
@@
When $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$, Theorem 5.2 shows how to compute $S$ given $R$ and $Q$, and Theorem 5.7 tells us that any two of $R,Q,S$ determines the third. With duality, it is straightforward, albeit a bit messy, to find explicit rod set formulas for $Q$ in terms of $R$ and $S$ and for $R$ in terms of $Q$ and $S$.
@@
For example, recall that we established the Fibonacci identities in Equations 3.2 and 3.3 using the third expansion tree in Figure 4 together with Theorem 5.5. We started with the infinite expansion
@@
Now we use Theorem 7.2 to dualize Equation 7.1:
@@
Theorem 3.3 implies the train counts for the rod set $1,3,5,7,\ldots]$ satisfy the Fibonacci recursion when $n>2$. The train counts for the rod set of odd numbers are the number of compositions, $\mathscr{O}(n)$, of $n$ into odd parts. Since $\mathscr{O}(1)=\mathscr{O}(2)=1$, the sequence $\mathscr{O}(n)$ is the Fibonacci sequence with these initial conditions, hence $\mathscr{O}(n)=\operatorname{F}(n-1,[1,2])$. For example, there are $\operatorname{F}(4,[1,2])=5$ rod trains of length $5$ built from odd length rods: $5,311,131,113,11111$.
@@
Since $R=[2,2,3]$ is an expansion of the Fibonacci rod set $[1,2]$ by the singleton rod set $[1]$, Corollary 5.6 implies $1,2]\xrightarrow{\hphantom{a}\hphantom{a}}S$ has discrepancies $3,6,9,\ldots]\cup[1,4,7,\ldots]$. The discrepancies are the rods with lengths $n\equiv 0,1\pmod{3}$, which leads to the following identities for the Fibonacci sequence $\operatorname{F}(n)=\operatorname{F}(n,[1,2])$:
@@ 8 Rod Sets and Binomial Coefficients
The two ways of writing this sum correspond to the dark diagonals in Figure 8. The first sum counts rods of length $b$ and the second rods of length $a$. Above the dark diagonal, the diagonal for $\operatorname{F}(n,[a,b])=0$ is empty (dashed in the figure) unless $n$ is a linear combination of $a$ and $b$. Below the dark diagonal the train counts are positive. In practice, it’s hard to draw the diagonals for large $a$. Hopkins [8, Appendix A] shows similar diagrams.
@@
Theorem 8.2 leads to some nice binomial identities. The train counts for $R=[1,1]$ are powers of two. The corresponding “diagonals” are the rows of the triangle, providing another proof of the familiar identity
@@ 9 Periodicity
Theorem 5.9 says that that when $n>4$, the net train counts $\operatorname{F}(n,[1,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{2}}])$ satisfy the recursion $\operatorname{F}(n)=\operatorname{F}(n-6)$ determined by rod set $[6]$.
@@
This is the special case of Theorem 5.9 when $S$ contains just a single positive rod. ∎
@@
Since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] expands via $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}$}] to $[2k]$, Corollary 5.6 implies $R$ expands to $[2k]$ via $Q\cup[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}]\cup Q{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{k}}$. ∎
@@
Theorem 9.2 says periodicity is equivalent to $R$ expanding to the one element rod set $[p]$, for which $\operatorname{\mathcal{C}}(x,[p])=x^{p}$. When $R$ is finite Theorem 6.3 then implies (1) and (5) are equivalent.
@@
We encourage the reader to check this result using Algorithm 5.8. Theorem 9.3 says we can stop when we’ve expanded to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{15}$}.
@@ 10 Expandability and Scaling
Since $\max Q=2$, Theorem 5.9, tells us that the $\operatorname{D}(n,R,[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{4}}^{4}])$ must vanish for $n\geq 3$. In particular, $\operatorname{F}(3)+4F(3-4)=\operatorname{F}(3)=0$ and $\operatorname{F}(4)+4F(0)=-4+4=0$.
@@
Theorem 5.9 implies $a=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>\max Q=a-\max R$. Thus for positive $i<\max R$
@@
To prove the converse, suppose $\operatorname{F}(a-i,R)=0$ for $i=1,2,\ldots,\max R-1$. Let $Q$ be the rod set that expands $R$ to the rod set $S$ with $\operatorname{F}(a,R)$ copies of $a$, with $a$ an antirod if $\operatorname{F}(a,R)$ is negative. The assumption on the net train counts of $R$ say that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $a-\max R<n<a$. Setting $\operatorname{C}(a,S)=\operatorname{F}(a,R)$ guarantees discrepancy 0 at $n=a$. Then Corollary 5.11 implies $Q$ is finite. ∎
@@
In Theorem 10.4 we saw that 1-expansions of $R$ correspond to sequences of 0’s in $\operatorname{F}(n,R)$. For 2-expandability the correspondence is to divisibility among net train counts. For example, we will see soon that the expansion above of $[1,2]$ to $8^{48},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{12}}^{7}]$ is equivalent to the fact that
@@
The Exchange Corollary 5.3 implies
@@
Suppose there is a finite expansion $R\xrightarrow{\hphantom{a}Q\hphantom{a}}S$ with $\operatorname{Shape}(S)=\langle a,b\rangle$. Theorem 5.9 implies $b=\max R+\max Q$, so the discrepancies measuring the ways the counts $\operatorname{F}(n,R)$ fail to satisfy the recursion for $S$ are 0 for $n>b-\max R$. In the range $b-\max R<n<b$, we have $\operatorname{F}(n-b,R)=0$.
@@
Let $Q$ be the rod set that expands $R$ to the $S$ with shape $\langle a,b\rangle$ and those net rod counts. It follows that the discrepancy $\operatorname{C}(n,Q)=\operatorname{D}(n,R,S)=0$ when $b-\max R<n\leq b$. Corollary 5.11 implies $Q$ is finite. ∎